Calculate the concentration (in molarity) of an NaOH solution if 25.0 mL of the solution is needed to neutralize 15.5 mL of a 0.305 M HCl solution.
Answer #1

From the equation of equivalance,

M(NaOH) ×V(NaOH) = M(HCl) × M(HCl)

M(NaOH) = M(HCl) × M(HCl) /  V(NaOH)

= (0.305×15.5 )/ 25.0= 0.1891